joemanner:

Use the matrix tool to solve the system of equations. Choose the correct ordered pair. 3x - 4y = -8 7x - 5y = 3

Ferrari:

Aight bet.

Ferrari:

Do you know how to solve matrix equations/expressions?

Ferrari:

Well, since you are offline, you can come back to the answer.

Ferrari:

First, add 4y to both sides; 3x - 4y+ 4y= -8 + 4y

Ferrari:

Next, you simplify that part. 3x = -8 + 4y

Ferrari:

Then, keeping the rule of whatever you do to one side you must do to the other, divide both sides of the equation by 3. \frac{3x}{3}=-\frac{8}{3}+\frac{4y}{3}

Ferrari:

Hm let me type the fraction out.

Ferrari:

\[\frac{ 3x }{ 3 } = -\frac{ 8 }{ 3 } + \frac{ 4y }{ 3 }\]

Ferrari:

Since we did this after adding 4y, we must simplify!

1 attachment
Ferrari:

Try doing the next one based off of what I explained to you for the first one.